# Ultrafast generation of coherent soft-shear phonons in halide perovskites via anisotropic photostriction

### v1

@@ Methods
Density functional theory calculations. The DFT calculations were performed using the WIEN2k package [55]. For the calculation of the elastic tensor components we use the IRelast package [56]. First, the structure is optimized using the PBEsol exchange-correlation functional. Next, for the few selected deformations the energy as a function of deformation amplitude is calculated and from the fit of the calculated energy, combinations of elastic tensor components are extracted.
@@
Ultrafast generation of coherent soft-shear phonons in halide perovskites via anisotropic photostriction
@@
Figure S1: (a) XRD powder pattern of the grained Cs2AgBiBr6 crystals measured at room temperature. The peaks can be assigned to a simple cubic crystal structure [41]. Peaks expected, but not resolved are marked with asterisks. (b) Simulated XRD powder pattern of the Cs2AgBiBr6 perovskite. The simulated data were generated with the Mercury CCDC software using the crystallographic data from [54].
@@
The tensor of elastic stiffness is defined [47] as
@@
The photoelastic (opto-elastic) tensor is defined [47] through the linear dependence of the components of the dielectric impermeabilities (inversed dielectric tensor) on the components of strain tensor:
@@
The tetragonal phase has the space group #87, class $4/m$, and the tensors are different. Both may be found in [47], § 1.1.4.10.6.5 (Note that there is an error in Table 1.6.7.1 of the same volume)
@@
It is worth to mention that, as explained in Refs. [51, 52], a change of the dielectric constant in low symmetry crystals is not fully captured by the tensor $p_{ij}$, but contains also an antisymmetric part which describes the change of the refraction index by twist stress. However, as explained in [53], these terms are zero if the refractive index is isotropic. According to the experimental data, the anisotropy of the refractive index in our case may be neglected and the photoelastic interaction may be described by the tensor (S5).
@@
Note also that the $\beta$-phase of perovskites has a slightly higher symmetry (space group #140, class $4/mmm$), where $p_{16}=p_{61}=p_{45}=0$ (we remind that there is an error in Ref. 47 in Table 1.6.7.1.).
Since the elastic stiffness tensor is not known for this material, we performed density functional theory (DFT) calculations to estimate the mode mixing in bulk material. The DFT calculations were done using the WIEN2k package [55]. For calculation of the elastic tensor components we use the IRelast package [56]. First, the structure is optimized using the PBEsol exchange-correlation functional (the optimized cubic lattice constant is 11.19 Å and in the tetragonal phase $a=7.84$, $c=11.32$). Next, for a few selected deformations, the energy as function of deformation amplitude is calculated and from the fit of the calculated energy the combinations of the elastic tensor components are extracted. The procedure gives the following set of constants in the cubic phase (constants are in GPa):
@@
$\left(\rho\omega^{2}\delta_{il}-c_{ijkl}k_{j}k_{k}\right)U_{l}({\bf Q},\omega)=\Phi_{i}({\bf Q},\omega)\,,$ (S14)
@@
$\Phi_{i}({\bf Q},\omega)=-n_{i}\alpha_{i}\rho_{X,S}\left[2\pi\delta(\omega)-\frac{2i}{\omega}\right]\frac{2n_{x}n_{y}n_{z}\zeta}{(n_{x}^{2}+k_{x}^{2})(n_{y}^{2}+k_{y}^{2})(n_{z}^{2}+k_{z}^{2})}.$ (S15)
@@
It can be demonstrated that the electric field of the scattered light can be written (see e.g. Refs. [50, 24]) in first order as

### v2

@@ Methods
Density functional theory calculations. The DFT calculations were performed using the WIEN2k package [Blaha2020]. For the calculation of the elastic tensor components we use the IRelast package [Jamal2018]. First, the structure is optimized using the PBEsol exchange-correlation functional. Next, for the few selected deformations the energy as a function of deformation amplitude is calculated and from the fit of the calculated energy, combinations of elastic tensor components are extracted.
@@
Efficient launch of shear phonons in photostrictive halide perovskites
@@
Figure S1: (a) XRD powder pattern of the grained Cs2AgBiBr6 crystals measured at room temperature. The peaks can be assigned to a simple cubic crystal structure [Adam-2016]. Peaks expected, but not resolved are marked with asterisks. (b) Simulated XRD powder pattern of the Cs2AgBiBr6 perovskite. The simulated data were generated with the Mercury CCDC software using the crystallographic data from [doi:10.1021/jacs.7b01629].
@@
The tensor of elastic stiffness is defined [ITCBD] as
@@
The photoelastic (opto-elastic) tensor is defined [ITCBD] through the linear dependence of the components of the dielectric impermeabilities (inversed dielectric tensor) on the components of strain tensor:
@@
The tetragonal phase has the space group #87, class $4/m$, and the tensors are different. Both may be found in [ITCBD], § 1.1.4.10.6.5 (Note that there is an error in Table 1.6.7.1 of the same volume)
@@
It is worth to mention that, as explained in Refs. [Nelson70, Anastassakis74], a change of the dielectric constant in low symmetry crystals is not fully captured by the tensor $p_{ij}$, but contains also an antisymmetric part which describes the change of the refraction index by twist stress. However, as explained in [Vacher72], these terms are zero if the refractive index is isotropic. According to the experimental data, the anisotropy of the refractive index in our case may be neglected and the photoelastic interaction may be described by the tensor (S5).
@@
Note also that the $\beta$-phase of perovskites has a slightly higher symmetry (space group #140, class $4/mmm$), where $p_{16}=p_{61}=p_{45}=0$ (we remind that there is an error in Ref. ITCBD in Table 1.6.7.1.).
Since the elastic stiffness tensor is not known for this material, we performed density functional theory (DFT) calculations to estimate the mode mixing in bulk material. The DFT calculations were done using the WIEN2k package [Blaha2020]. For calculation of the elastic tensor components we use the IRelast package [Jamal2018]. First, the structure is optimized using the PBEsol exchange-correlation functional (the optimized cubic lattice constant is 11.19 Å and in the tetragonal phase $a=7.84$, $c=11.32$). Next, for a few selected deformations, the energy as function of deformation amplitude is calculated and from the fit of the calculated energy the combinations of the elastic tensor components are extracted. The procedure gives the following set of constants in the cubic phase (constants are in GPa):
@@
$\left(\rho\omega^{2}\delta_{il}-c_{ijkl}k_{j}k_{k}\right)U_{l}({\bf k},\omega)=\Phi_{i}({\bf k},\omega)\,,$ (S14)
@@
$\Phi_{i}({\bf k},\omega)=-n_{i}\alpha_{i}\rho_{X,S}\left[2\pi\delta(\omega)-\frac{2i}{\omega}\right]\frac{2n_{x}n_{y}n_{z}\zeta}{(n_{x}^{2}+k_{x}^{2})(n_{y}^{2}+k_{y}^{2})(n_{z}^{2}+k_{z}^{2})}.$ (S15)
@@
It can be demonstrated that the electric field of the scattered light can be written (see e.g. Refs. [Subbaswamy1978, Matsuda2008]) in first order as